OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line Y= 3/2x +4 and passes through the point (−6, 3).

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

First, identify the slope of the first equation. From the formula y=mx+b, you can see that the slope of this graph is 3/2. For a perpendicular line, the slope will be the negative reciprocal, so -2/3. Plug this into the equation for a line to get y=(-2/3)x+b. In order to find out what b is, plug the point you are given into this equation. 3=(-2/3)(-6)+b 3=4+b b=3-4 b=-1 Plug that back into the equation, and you get the final answer: y=(-2/3)x-1